Zeltron is an all-equity firm with 130 million shares outstanding, which are currently trading for $11.57 per share. A month ago, Zelatron announced it will change its capital structure by borrowing $211 million in short-term debt borrowing $244 million in long term debt and issuing $213 million of preferred stock. The 688 million raised by these issues plus another $23 million in cash that Zeltron already has, will be used to repurchase existing shares of stock. The transaction is scheduled to occur today. Assume perfect capital markets.

A. What is the market value balance sheet of Zeltron?

1. Before the transaction

2. After the new securities are issues but the share repurchase

3. After the share repurchase

B. At the conclusion of this transaction how many shares outstanding will Zeltron have, and what will the value of those shares be?

Sorry it is 668
Answer #1

Answer :

A)Market value Balance sheet

1) Before Transaction

Equity =Market value per share *Shares outstanding

Liability

Equity=11.57*130 million =$1504.1 million

Asset

Cash=$23 million

2) After new security and before repurchase of share

Liability

Short term Debt=$211 million

Long term Debt= $244 million

Prefferd Stock= $213 million

Equity =$1504.1 million

Asset

Cash =$23 million

3) after repurchase of shares

Since total amount for shares repurchase = 211+244+213+23=$691 millions

So total shares repurchased = 691/11.57=59.72 millions

So market value of remaining shares =1504.1-691=$813.1 millions

Liability

Short term debt =$211 million

Long term debt =$244 million

Preferred shares=$213 million

Equity =$813.1 million

B)

Shares outstanding =130-59.72=$70.28 million

Value of these shares =$813.1 million